# Eigenstate thermalization to non-monotonic distributions in strongly-interacting chaotic lattice gases

### v2

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
@@
ETH means that the eigenstate to eigenstate fluctuations of expectation values within any chaotic microcanonical shell are suppressed. In certain situations [11, 12], the fluctuation variance is inversely-proportional to the number of principal components (NPC) $\mathcal{N}_{\mathrm{PC}}$ — the estimate of the number of integrable system eigenstates comprising the non-integrable one. Thus, ETH typically implies large NPC, but can be practically attained when NPC is substantially smaller than the dimension $\mathcal{N}_{\mathrm{HS}}$ of the Hilbert space (which can be also constrained due to possible conservation laws). NPC approaches a large fraction (limited by $1/3$ for time-reversible systems [13] or $1/2$ for time-irreversible ones [14]) of $\mathcal{N}_{\mathrm{HS}}$ only in the regime of quantum ergodicity [15].
@@
2 The lattice models
[…]
@@ 2 The lattice models
The second model is a one-dimensional (1D) Bose-Hubbard (BH) chain with $N$ spinless bosons in $L$ sites, with on-site interactions of strength $V$ and hard wall boundaries (see Appendix B). Parity symmetry is broken by adding a random disorder/bias potential of order $\leq 0.05$. The resulting Hilbert space dimension for the bosonic system is $\mathcal{N}_{\mathrm{HS}}=(N+L-1)!/(N!(L-1)!)$. The system with $N=10$ particles in $L=8$ sites, considered here, has $\mathcal{N}_{\mathrm{HS}}=19448$.
@@
Analyzing the chaotic system properties, we have to compare them to ones of the closest integrable system. For this purpose, we use corresponding systems of non-interacting particles. Their symmetric or anti-symmetric many-body eigenfunctions — the orbital Fock states $\left|n\right\rangle=\left|n_{1}\ldots n_{L}\right\rangle$ — have the eigenenergies $E_{n}=\sum_{k}n_{k}\varepsilon_{k}$. Here $n_{k}$ are occupations of the 1B orbitals, labeled in increasing order of their eigenenergies $\varepsilon_{k}$. Subtractions the average expectation values of interactions from the interacting particle Hamiltonians (see Appendices A and B) leads to a substantial overlap of the non-interacting and interacting spectra $\{E_{n}\}$ and $\{E_{\alpha}\}$
@@
Figure 1: 1B orbital occupations for (a) the FH model with the interaction strengths $V=0$ (dashed lines), $V=1$ (pluses), and $V=10$ (solid lines), averaged over microcanonical shells with the mean energies $-6$ (black), 00 (green), and $6$ (red) and (b) the 1D BH model with the interaction strengths $V=0$ (dashed lines), $V=0.3$(pluses), and $V=3$ (solid lines), averaged over microcanonical shells with the mean energies $-1.2$ (black), $-0.17$ (green), and $1.2$ (red).
3 Exact diagonalization results
@@
For sufficiently strong interaction, both models become chaotic. For the FH model at $V=1$ the ratio of two consecutive level spacings [29], averaged over the energy spectrum (see Appendix C) increases to $\left\langle r\right\rangle\approx 0.525$ (cf. $\left\langle r\right\rangle\approx 0.536$ [30] for the Wigner-Dyson ensemble of Gaussian orthogonal matrices, describing completely-chaotic systems). The chaotic behavior is confirmed also by suppression of eigenstate-to-eigenstate fluctuations of the observable expectation values (see Appendix C). Their variances are reduced by two orders of magnitude. Another criterion of chaoticity, NPC, increased to $5\times 10^{2}$ (see Appendix C). The chaoticity of the BH model is determined by the value of $VN$. For the BH model at $VN=3.0$ ($V=0.3$) we have $\left\langle r\right\rangle\approx 0.529$, fluctuation variances are reduced by two orders of magnitude, and NPC is increased to $8\times 10^{2}$. For these interaction strengths, the microcanonical distributions of the 1B orbital occupations  [^1] The orbital occupations can be directly observed in experiments. In addition, all single-particle parameters can be expressed in terms of them. for interacting and non-interacting particles are very close (see Fig. 1). This is in line with [10], as subtracting the average expectation values of interactions from the interacting particle Hamiltonians is equivalent to the energy shift used there. We notice, that both interacting and non-interacting distributions are different from the FD and BE distribution due to the small system size [28]. Due to macroscopic self-trapping, the BH model becomes integrable again at large $V$ where the site populations become effective integrals of motion. Thus, the chaoticity parameter reduces to $\left\langle r\right\rangle\approx 0.41$ at $V=10$ (cf $\left\langle r\right\rangle\approx 0.386$ for integrable systems).
@@ 4 Non-monotonic distributions
$\overline{O}(E)=\frac{\Delta_{n}}{\Delta_{\mathrm{MC}}}\sum_{n\in\mathrm{MC}(E)}O_{n},$
where $\Delta_{n}$ is the average distance between neighboring $E_{n}$ in the vicinity of $E$. Further, as $\Gamma$ substantially exceeds $\Delta_{n}$, approximating summation in (2) by integration, we get
$\overline{\left\langle\alpha\left|\hat{O}\right|\alpha\right\rangle}\approx\int_{E_{\mathrm{min}}}^{E_{\mathrm{max}}}\frac{dE^{\prime}}{\Delta_{n}}\Delta_{\alpha}W(E,E^{\prime})\overline{O}(E^{\prime}),$ (4)
@@
Consider a particular case of the orbital occupation operator $\hat{N}_{k}\left|n\right\rangle=\sum_{j=1}^{N}n_{k_{j}}\left|n\right\rangle$. The shape of the microcanonical distribution of the orbital occupations $\overline{N_{k}}(E)$ alters with the mean shell energy (see Fig. 1). If $W(E,E^{\prime})$ vanishes when $|E-E^{\prime}|>\Gamma$ and $\Gamma$ is small with respect to the energy scale on which the microcanonical distribution varies, we have $\overline{\left\langle\alpha\left|\hat{N}_{k}\right|\alpha\right\rangle}\equiv\overline{N_{k}^{int}(}E)\approx\overline{N_{k}}(E)$, thus justifying the equivalence between the occupation statistics of the interacting and non-interacting systems. However, if $\Gamma$ exceeds this scale, the interacting-system’s occupation distribution $\overline{N_{k}^{int}(}E)$ can mix non-interacting distributions $\overline{N_{k}}(E)$ of different shape and be different from any individual non-interacting microcanonical distribution $\overline{N_{k}}(E)$. In Fig. 1, mixing of increasing and decreasing $\overline{N_{k}}(E)$ leads to near uniform $\overline{N_{k}^{int}(}E)$
The exact diagonalization method is applicable only to small numbers of particles and lattice sites when the microcanonical distribution of the orbital occupations $\overline{N_{k}}(E)$ is different from the FD and BE distributions [28]. However, for large numbers of particles the microcanonical occupation of the orbitals $\overline{N_{k}}(E)$ are precisely given by the FD or BE distributions
@@
While finding the exact strength-function by direct diagonalization is not possible for large systems, in the case of strong interactions, it can be approximated by the Gaussian shape (see [13])
$W(E,E_{n})\approx C(E)\frac{\Delta_{n}}{\Delta_{\alpha}}\exp(-(E-E_{n})^{2}/\Gamma^{2}).$ (6)
where $\Delta_{n}$ is taken in the vicinity of $E_{n}$ and the normalization factor $C(E)$ is determined by
@@
The Gaussian shape approximates the LDOS with good accuracy even for systems of small size (see Appendix D). It should be stressed, that the agreement can be provided by the factor $\Delta_{n}$. The resulting distributions, calculated with Eqs. (4), (5), (6), and (7) depend on scaled widths $\tilde{\Gamma}=\Gamma/L$. In addition to the 2D FH and 1D BH model, treated above using exact diagonalization, we consider also the 2D BH model with the same 1B Hamiltonian as the 2D FH one (see Appendix A). Figure 2 shows the obtained distributions for the Gaussian width $\Gamma$ (increasing with the interaction strength $V$) covering both eigenstates corresponding to positive and negative temperature, or, respectively, to the decreasing and increasing FD or BE distributions. The resulting distributions for the interacting system have clearly pronounced minima. This is provided by the positive second derivative $d^{2}\overline{N_{k}}(E)/d\varepsilon^{2}_{k}>0$ for the FD and BE distributions (5) when $(\varepsilon_{k}-\mu)/T>0$. By contrast, for small numbers of particles the minimum does not appear due to a maximum of $\overline{N_{k}}(E)$ ($d^{2}\overline{N_{k}}(E)/d\varepsilon^{2}_{k}<0$) with $E\approx 0$ (see Fig. 1), and the maximum vanishes in $\overline{N_{k}^{int}(}E)$ due to strong interactions.
5 Discussion
In the MBL literature, the proportionality of NPC to Hilbert space dimension is used as an attribute of delocalization, distinguishing extended eigenstates from localized eigenstates. This property was reported for eigenstates in Heisenberg [31] and XXZ [32] spin chains, the Bose-Hubbard model [33], and the Jaynes-Cummings-Hubbard system [34]. Then, the ratio $\tilde{\Gamma}/\tilde{E}\sim\tilde{N}\Delta\varepsilon\mathcal{N}_{\mathrm{PC}}/(\mathcal{N}_{\mathrm{HS}}\tilde{E})$ should remain unchanged for extended eigenstates in the thermodynamic limit $N\rightarrow\infty$, while $\tilde{N}=\mathrm{const}$ and $\tilde{E}=\mathrm{const}$. As a result, the distribution deviations from the FD and BE ones may survive in the thermodynamic limit.
@@ 5 Discussion
We reemphasize that the above deviations from the FD or BE distributions are quite different from those observed in [10]. The distributions presented in Fig. 2 therein are monotonic and agree perfectly with the BE distributions for the temperatures corresponding to the dressed energies. The dressed energy, given by Eq. (12) in [10], is shifted by the average expectation value of the interactions. As noted earlier, this prescription is implicit in all our calculations, as the average expectation values of the interactions are subtracted from the interacting-system eigenenergies (see Appendices A and B). The deviations we observe are more profound than this simple energy shift. Our main point is that the interactions mix different microcanonical shells of the non-interacting system, so that the microcanonical occupation means over the interacting system’s energy shell do not match any of the corresponding microcanonical means over non-interacting shells. This voids all energy shift prescriptions, including that of [10]. The non-monotonic distributions obtained here are a result of this mixing of shells with positive- and negative temperature and can not be reduced to a change of temperature.
6 Conclusion
@@
Acknowledgements
AV acknowledges support from the NSF through a grant for ITAMP at Harvard University.
@@ Appendix A Two-dimensional lattice models
$\hat{H}_{F}$ $=-\sum_{l_{x}=1}^{L_{x}}\sum_{l_{y}=1}^{L_{y}}\sum_{\delta_{x}=-1}^{1}\sum_{\delta_{y}=-1}^{1}\left(1-\delta_{\delta_{x}0}\delta_{\delta_{y}0}\right)\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}+\delta_{x}l_{y}+\delta_{y}}+\hat{V}_{F}-\bar{V},$ (A.1)
$\hat{V}_{F}$ $=V\sum_{l_{x}=1}^{L_{x}}\sum_{l_{y}=1}^{L_{y}}\left(\sum_{\delta_{x}=\pm 1}\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}+\delta_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}l_{y}}\hat{a}_{l_{x}+\delta_{x}l_{y}}+\sum_{\delta_{y}=\pm 1}\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}l_{y}+\delta_{y}}^{\dagger}\hat{a}_{l_{x}l_{y}}\hat{a}_{l_{x}l_{y}+\delta_{y}}\right),$ (A.2)
@@
$p_{x}=\frac{2\pi m_{x}+\chi_{x}}{L_{x}}\quad(1\leq m_{x}\leq L_{x}),\quad p_{y}=\frac{2\pi m_{y}+\chi_{y}}{L_{y}}\quad(1\leq m_{y}\leq L_{y}),$ (A.3)
@@
[…]
@@
$\left\langle n\left|\hat{V}_{F}\right|n\right\rangle=\frac{4V}{L}\sum_{k<k^{\prime}}n_{k}n_{k}^{\prime}\left(\sin^{2}\frac{p_{x}(k)-p_{x}(k^{\prime})}{2}+\sin^{2}\frac{p_{y}(k)-p_{y}(k^{\prime})}{2}\right).$ (A.7)

### v4

@@ Abstract
I Introduction
@@
The eigenstates of the non-integrable system are superpositions of the integrable-system eigenstates. The number of principal components (NPC) $\mathcal{N}_{\mathrm{PC}}$ estimates (for each exact eigenstate) the number of contributing integrable-system eigenstates (see App. C). ETH means that the eigenstate to eigenstate fluctuations of expectation values within any chaotic microcanonical shell are suppressed. In certain situations [11, 12], the fluctuation variance is inversely-proportional to $\mathcal{N}_{\mathrm{PC}}$. Thus, ETH typically implies large NPC, but can be practically attained when NPC is substantially smaller than the dimension $\mathcal{N}_{\mathrm{HS}}$ of the Hilbert space (which can be also constrained due to possible conservation laws). NPC approaches a large fraction (limited by $1/3$ for time-reversible systems [13] or $1/2$ for time-irreversible ones [14]) of $\mathcal{N}_{\mathrm{HS}}$ only in the regime of quantum ergodicity [15].
@@
II The lattice models
We find eigenstates of two lattice models by direct numerical diagonalization, allowed up to the Hilbert space dimension $\mathcal{N}_{\mathrm{HS}}\lesssim 2\times 10^{4}$. Throughout this manuscript, all energies are measured in units of the lattice hopping energy. In the first, two-dimensional (2D) Fermi-Hubbard (FH) model, $N$ spin-polarized fermions have nearest-neighbor interactions with the strengths $V$ (see App. A). This model includes hoppings with simultaneous change of $l_{x}$ and $l_{y}$, which label sites of the $L_{x}\times L_{y}$ lattice. This, together with twisted-periodic boundary conditions, allow us to remove degeneracies of the many-body non-interacting particle eigenstates. The total number of one-body (1B) states in this model is $L=L_{x}L_{y}$. Due to the spatial homogeneity of this model, we consider separately each sector with the given total momentum which contains $\mathcal{N}_{\mathrm{HS}}\approx(L-1)!/(N!(L-N)!)$ eigenstates. The results below are obtained for $N=6$ particles in the $6\times 5$ lattice ($L=30$) and the total momentum $x$ and $y$ components $3$ and $2$, respectively. In this case, $\mathcal{N}_{\mathrm{HS}}=19811$.
@@ II The lattice models
The second model is a one-dimensional (1D) Bose-Hubbard (BH) chain with $N$ spinless bosons in $L$ sites, with on-site interactions of strength $V$ and hard wall boundaries (see App. B). Parity symmetry is broken by adding a random disorder/bias potential of order $\leq 0.05$ (see also [29]). The resulting Hilbert space dimension for the bosonic system is $\mathcal{N}_{\mathrm{HS}}=(N+L-1)!/(N!(L-1)!)$. The system with $N=10$ particles in $L=8$ sites, considered here, has $\mathcal{N}_{\mathrm{HS}}=19448$.
@@
Analyzing the chaotic system properties, we have to compare them to ones of the closest integrable system. For this purpose, we use corresponding systems of non-interacting particles. Their symmetric or anti-symmetric many-body eigenfunctions — the orbital Fock states $\left|n\right\rangle=\left|n_{1}\ldots n_{L}\right\rangle$ — have the eigenenergies $E_{n}=\sum_{k}n_{k}\varepsilon_{k}$. Here $n_{k}$ are occupations of the 1B orbitals, labeled in increasing order of their eigenenergies $\varepsilon_{k}$. Subtractions the average expectation values of interactions from the interacting particle Hamiltonians (see Apps. A and B) leads to a substantial overlap of the non-interacting and interacting spectra $\{E_{n}\}$ and $\{E_{\alpha}\}$ .
@@
Figure 1: 1B orbital occupations for (a) the FH model with the interaction strengths $V=0$ (dashed lines), $V=1$ (pluses), and $V=10$ (solid lines), averaged over microcanonical shells with the mean energies 6-6 (black), 0 (green), and 66 (red) and (b) the 1D BH model with the interaction strengths $V=0$ (dashed lines), V=0.3V=0.3(pluses), and $V=3$ (solid lines), averaged over microcanonical shells with the mean energies 1.2-1.2 (black), 0.17-0.17 (green), and $1.2$ (red).
III Exact diagonalization results
@@
For sufficiently strong interaction, both models become chaotic. For the FH model at $V=1$ the ratio of two consecutive level spacings [30], averaged over the energy spectrum (see App. C) increases to $\left\langle r\right\rangle\approx 0.525$ (cf. $\left\langle r\right\rangle\approx 0.536$ [31] for the Wigner-Dyson ensemble of Gaussian orthogonal matrices, describing completely-chaotic systems). The chaotic behavior is confirmed also by suppression of eigenstate-to-eigenstate fluctuations of the observable expectation values (see App. C). Their variances are reduced by two orders of magnitude. Another criterion of chaoticity, NPC, increased to $5\times 10^{2}$ (see App. C). The chaoticity of the BH model is determined by the value of $VN$. For the BH model at $VN=3.0$ ($V=0.3$) we have $\left\langle r\right\rangle\approx 0.529$, fluctuation variances are reduced by two orders of magnitude, and NPC is increased to $8\times 10^{2}$. For these interaction strengths, the microcanonical distributions of the 1B orbital occupations  [^1] The orbital occupations can be directly observed in experiments. In addition, all single-particle parameters can be expressed in terms of them. for interacting and non-interacting particles are very close (see Fig. 1). This is in line with [10], as subtracting the average expectation values of interactions from the interacting particle Hamiltonians is equivalent to the energy shift used there. We notice, that both interacting and non-interacting distributions are different from the FD and BE distribution due to the small system size [28]. Due to macroscopic self-trapping, the BH model becomes integrable again at large $V$ where the site populations become effective integrals of motion. Thus, the chaoticity parameter reduces to $\left\langle r\right\rangle\approx 0.41$ at $V=10$ (cf $\left\langle r\right\rangle\approx 0.386$ for integrable systems).
@@ IV Non-monotonic distributions
$\overline{O}(E)=\frac{\Delta_{n}(E)}{\Delta_{\mathrm{MC}}}\sum_{n\in\mathrm{MC}(E)}O_{n},$ (4)
where $\Delta_{n}(E)$ is the average distance between neighboring $E_{n}$ in the vicinity of $E$. Further, as $\Gamma$ substantially exceeds $\Delta_{n}$, approximating summation in (2) by integration, we get
$\overline{\left\langle\alpha\left|\hat{O}\right|\alpha\right\rangle}\approx\int_{E_{\mathrm{min}}}^{E_{\mathrm{max}}}\frac{dE^{\prime}}{\Delta_{n}(E^{\prime})}\Delta_{\alpha}(E)W(E,E^{\prime})\overline{O}(E^{\prime}),$ (5)
@@
Consider a particular case of the 1B orbital occupation operator $\hat{N}_{k}\left|n\right\rangle=n_{k}\left|n\right\rangle$, where $k$ labels the 1B orbitals in increasing order of their eigenenergies $\varepsilon_{k}$ (see Sec. II and Apps. A and B). The microcanonical distribution of the orbital occupations for non-interacting particles $\overline{N_{k}}(E)$ is given by Eq. (4). The shape of $\overline{N_{k}}(E)$ depends on the mean shell energy $E$ (see Fig. 1). If $W(E,E^{\prime})$ vanishes when $|E-E^{\prime}|>\Gamma$ and $\Gamma$ is small with respect to the energy scale on which the microcanonical distribution $\overline{N_{k}}(E)$ varies, we can approximate the microcanonical distribution for interacting particles $\overline{N_{k}^{int}}(E)\equiv\overline{\left\langle\alpha\left|\hat{N}_{k}\right|\alpha\right\rangle}$ [see Eq. (5))] as $\overline{N_{k}^{int}}(E)\approx\overline{N_{k}}(E)$, thus justifying the equivalence between the occupation statistics of the interacting and non-interacting systems. However, if $\Gamma$ exceeds this scale, the interacting-system’s occupation distribution $\overline{N_{k}^{int}}(E)$ can mix non-interacting distributions $\overline{N_{k}}(E)$ of different shape and be different from any individual non-interacting microcanonical distribution $\overline{N_{k}}(E)$. In Fig. 1, mixing of increasing and decreasing $\overline{N_{k}}(E)$ leads to near uniform $\overline{N_{k}^{int}}(E)$ .
The exact diagonalization method is applicable only to small numbers of particles and lattice sites when the microcanonical distribution of the orbital occupations $\overline{N_{k}}(E)$ is different from the FD and BE distributions [28]. However, for large numbers of non-interacting particles the microcanonical and canonical thermodynamic means become equivalent (although the fluctuations in different ensembles can be nonequivalent even in the thermodynamic limit [33, 34]). Then, the microcanonical occupations of the orbitals $\overline{N_{k}}(E)$ are precisely given by the FD or BE distributions
@@
While finding the exact LDOS by direct diagonalization is not possible for large systems, in the case of strong interactions, it can be approximated by the Gaussian shape (see [13])
$W(E,E_{n})\approx C(E)\frac{\Delta_{n}(E_{n})}{\Delta_{\alpha}(E)}\exp(-(E-E_{n})^{2}/\Gamma^{2}).$ (7)
where $\Delta_{n}(E_{n})$ is taken in the vicinity of $E_{n}$ and the normalization factor $C(E)$ is determined by
$1/C(E)=\int_{E_{\mathrm{min}}}^{E_{\mathrm{max}}}\exp(-(E-E^{\prime})^{2}/\Gamma^{2})dE^{\prime}.$ (8)
The Gaussian shape (7) approximates the LDOS with good accuracy even for systems of small size (see App. D). It should be stressed, that the agreement can be provided by the factor $\Delta_{n}(E_{n})$ even if the spectrum is substantially inhomogeneous. The resulting distributions, calculated with Eqs. (5), (6), (7), and (8) depend on scaled widths $\tilde{\Gamma}=\Gamma/L$. In addition to the 2D FH and 1D BH model, treated above using exact diagonalization, we consider also the 2D BH model with the same 1B Hamiltonian as the 2D FH one (see App. A). Figure 2 shows the obtained distributions for the Gaussian width $\Gamma$ (increasing with the interaction strength $V$) covering both eigenstates corresponding to positive and negative temperature, or, respectively, to the decreasing and increasing FD or BE distributions. The resulting distributions for the interacting system have clearly pronounced minima. This is provided by the positive second derivative $d^{2}\overline{N_{k}}(E)/d\varepsilon^{2}_{k}>0$ for the FD and BE distributions (6) when $(\varepsilon_{k}-\mu)/T>0$. In contrast, for the few-mode systems of Fig. 1, the minimum does not appear due to a maximum of $\overline{N_{k}}(E)$ ($d^{2}\overline{N_{k}}(E)/d\varepsilon^{2}_{k}<0$) with $E\approx 0$, and the maximum vanishes in $\overline{N_{k}^{int}(}E)$ due to strong interactions.
V Discussion
It should be realized that Fig. 1 and Fig. 2 show the same effect, namely the mixing of occupation distributions obtained from different microcanonical shells of the non-interacting system, due to a broadened LDOS. However, whereas in Fig. 2 the mean natural-mode occupations within microcanonical shells of the non-interacting system take the ubiquitous FD or BE forms (6), the corresponding distributions for few-mode mesoscopic system that can be numerically diagonalized as in Fig. 1, have non-universal and sometimes non-monotonic structure in themselves [28]. The resulting strong-interaction occupation distribution is hence qualitatively different in the two cases. While numerical complexity precludes direct diagonalization for large systems, the non-monotonic distributions in Fig. 2 are the inevitable result of the well-known FD or BE noninteracting shell distributions and the Gaussian lineshape of the LDOS (7) that is confirmed in numerical calculations [13] and extends also to smaller systems (see App.D). Moreover, even this particular lineshape is not necessary, as non-monotonic distributions eventually appear for LDOS of any shape as long as it is sufficiently broadened by strong interactions.
@@ V Discussion
We reemphasize that the above deviations from the FD or BE distributions are quite different from those observed in [10]. The distributions presented in Fig. 2 therein are monotonic and agree perfectly with the BE distributions for the temperatures corresponding to the dressed energies. The dressed energy, given by Eq. (12) in [10], is shifted by the average expectation value of the interactions. As noted earlier, this prescription is implicit in all our calculations, as the average expectation values of the interactions are subtracted from the interacting-system eigenenergies (see Apps. A and B). The deviations we observe are more profound than this simple energy shift. Our main point is that the interactions mix different microcanonical shells of the non-interacting system, so that the microcanonical occupation means over the interacting system’s energy shell do not match any of the corresponding microcanonical means over non-interacting shells. This voids all energy shift prescriptions, including that of [10]. The non-monotonic distributions obtained here are a result of this mixing of shells with positive- and negative temperature and can not be reduced to a change of temperature. Mixing of different shells was already considered in classical superstatistics [40] and thermodynamics of small systems [41], but with no relation to quantum eigenstates and non-monotonic distributions.
VI Conclusion
@@
Acknowledgement
VY and AV acknowledge support from the NSF through a grant for ITAMP at Harvard University.
@@ Appendix A Two-dimensional lattice models
$\hat{H}_{F}=$ $-\sum_{l_{x}=1}^{L_{x}}\sum_{l_{y}=1}^{L_{y}}\sum_{\delta_{x}=-1}^{1}\sum_{\delta_{y}=-1}^{1}\left(1-\delta_{\delta_{x}0}\delta_{\delta_{y}0}\right)\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}+\delta_{x}l_{y}+\delta_{y}}$
$+\hat{V}_{F}-\bar{V},$ (A.1)
$\hat{V}_{F}=$ $V\sum_{l_{x}=1}^{L_{x}}\sum_{l_{y}=1}^{L_{y}}\Bigl(\sum_{\delta_{x}=\pm 1}\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}+\delta_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}l_{y}}\hat{a}_{l_{x}+\delta_{x}l_{y}}$
$+\sum_{\delta_{y}=\pm 1}\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}l_{y}+\delta_{y}}^{\dagger}\hat{a}_{l_{x}l_{y}}\hat{a}_{l_{x}l_{y}+\delta_{y}}\Bigr),$ (A.2)
@@
$p_{x}=\frac{2\pi m_{x}+\chi_{x}}{L_{x}}\quad(1\leq m_{x}\leq L_{x})$
$p_{y}=\frac{2\pi m_{y}+\chi_{y}}{L_{y}}\quad(1\leq m_{y}\leq L_{y}),$
@@
$\frac{k(\varepsilon)}{L}\approx\frac{1}{(2\pi)^{2}}\int_{0}^{2\pi}dp_{x}\int_{0}^{2\pi}dp_{y}\vartheta(\varepsilon-\varepsilon_{2D}(p_{x},p_{y})),$ (A.5)
@@
$\left\langle n\left|\hat{V}_{F}\right|n\right\rangle=\frac{4V}{L}\sum_{k<k^{\prime}}$ $n_{k}n_{k}^{\prime}\Bigl(\sin^{2}\frac{p_{x}(k)-p_{x}(k^{\prime})}{2}$
$+\sin^{2}\frac{p_{y}(k)-p_{y}(k^{\prime})}{2}\Bigr).$ (A.7)